a shipment of computer monitors, some weighing 25lb and the others weighing 40 lb each, has a total weight of 680 lb. If there are 20 monitors altogether, how many weigh 40 lb?
Answer by checkley79(3341)

25x+40y=680
x+y=20
x=20-y
25(20-y)+40y=680
500-25y+40y=680
15y=680-500
15y=180
y=180/15
y=12 40 lb. monitors.
x+12=20
x=20-12
x=8 25 lb. monitors.
Proof:
25*8+40*12=680
200+480=680
680=680